# Empirical sandwich variance estimator for iterated conditional expectation g-computation

### v2

@@ Abstract
Iterated conditional expectation (ICE) g-computation is an estimation approach for addressing time-varying confounding for both longitudinal and time-to-event data. Unlike other g-computation implementations, ICE avoids the need to specify models for each time-varying covariate. For variance estimation, previous work has suggested the bootstrap. However, bootstrapping can be computationally intense and sensitive to the number of resamples used. Here, we present ICE g-computation as a set of stacked estimating equations. Therefore, the variance for the ICE g-computation estimator can be consistently estimated using the empirical sandwich variance estimator. Performance of the variance estimator was evaluated empirically with a simulation study. The proposed approach is also demonstrated with an illustrative example on the effect of cigarette smoking on the prevalence of hypertension. In the simulation study, the empirical sandwich variance estimator appropriately estimated the variance. When comparing runtimes between the sandwich variance estimator and the bootstrap for the applied example, the sandwich estimator was substantially faster, even when bootstraps were run in parallel. The empirical sandwich variance estimator is a viable option for variance estimation with ICE g-computation.
@@ 1 Introduction
To consistently estimate the variance for ICE g-computation, previous work has suggested the nonparametric bootstrap [7, 8, 10]. However, the bootstrap is computationally demanding, as it requires repeating the analysis using resamples of the data [11, 12]. Further, the bootstrap requires the analyst to specify the number of iterations, where a larger number of iterations is expected to reduce Monte Carlo error but will increase computation time [11, 13]. This computational complexity can limit the scenarios considered by researchers in practical applications (e.g., exploring alternative treatment plans, varying functional form specifications, sensitivity analyses). The computational complexity also makes simulation experiments difficult. For example, some simulation studies forgo estimation of the variance by bootstrap in each iteration [8, 9]. While other simulation studies have used the bootstrap [10], the computational burden may limit the sample sizes, number of iterations, or scenarios considered. An alternative variance estimator is based on the influence curve, but this estimator is not consistent for ICE g-computation [7].
Here, the ICE g-computation estimator is expressed as an M-estimator, i.e., a set of unbiased estimating equations [14], which allows the asymptotic variance of the ICE estimator to be consistently estimated using the empirical sandwich variance estimator [8]. The empirical sandwich variance estimator has also been used for variance estimation with g-computation in settings without time-varying confounding [15, 16, 17, 18]. Here, we illustrate the use of M-estimation for ICE g-computation in settings with time-varying confounding. The proposed approach has the usual benefits of ICE g-computation, in addition to providing a statistically consistent variance estimator that is more computationally efficient than the nonparametric bootstrap. This approach also allows one to easily estimate the variance for transformations of parameters and incorporate additional nuisance models by stacking estimating equations together. Finally, the ICE g-computation estimator can be readily implemented using existing software for general M-estimators [19, 20].
The structure of the paper is as follows. In section 2, the data and a sufficient set of identification assumptions for longitudinal data structures are reviewed. Section 3 reviews g-computation estimators in the setting of repeated measures and presents ICE g-computation as stacked estimating equations. The proposed M-estimator is assessed through a simulation study in section 4. In section 5, the proposed ICE g-computation procedure is demonstrated in an illustrative example of estimating the effect of cigarette smoking on prevalent hypertension. Finally, section 6 summarizes the key results and notes how an estimating equation approach for causal effect estimation with longitudinal data can be expanded upon in future work.
@@ 2 Observed data and identification
Let $k\in\{1,...,\tau\}$ index discrete follow-up times. The potential outcome at $\tau$ for unit $i$ under the treatment plan (i.e., a defined sequence of treatments) $\bar{a}_{\tau-1}^{*}=(a_{0}^{*},a_{1}^{*},...,a_{\tau-1}^{*})$ is denoted by $Y_{i,\tau}(\bar{a}_{\tau-1}^{*})$. Only deterministic plans of binary treatments (i.e., units are assigned to specific treatments, as opposed to probabilities of treatments) are considered hereafter. A simple example is always treat, where $\bar{a}_{\tau-1}^{*}=(1,1,...,1)$. The parameter of interest is the mean potential outcome at $\tau$ under a given plan (i.e., the joint effect of treatment), ${\color[rgb]{0,0,0}\mu_{\tau}(\bar{a}_{\tau-1}^{*})}=\mathbb{E}[Y_{i,\tau}(\bar{a}_{\tau-1}^{*})]$, where $\mathbb{E}[\cdot]$ is the expected value function.
@@
To identify ${\color[rgb]{0,0,0}\mu_{\tau}(\bar{a}_{\tau-1}^{*})}$, we proceed following the assumptions provided in Table 1. Comprehensive discussion of these assumptions can be found elsewhere [4, 8, 22, 23, 24, 25, 26]. Briefly, causal consistency provides a connection between the observed outcomes and time-varying covariates, and the potential outcomes and covariates under the treatment plans. Causal consistency, as expressed here, implies both no interference between units and variations of treatment are irrelevant [27]. Treatment exchangeability stipulates that the treatment at time $k$ is independent of the potential outcomes conditional on the history of previous treatment and covariates. In other words, $L$ includes all the confounding variables. In the context of deterministic plans, treatment positivity states that there is a non-zero probability of following the plan up to $k$ conditional on having followed the plan up to $k-1$ and time-varying confounding history that occur. Censoring exchangeability specifies that loss to follow-up at time $k$ is non-informative conditional on $\bar{L}_{k},\bar{A}_{k},\bar{Y}_{k}$. Again, censoring positivity ensures that the expression for censoring exchangeability is well-defined. In addition to these assumptions, we further assume no measurement error. Following these additional assumptions, the parameter of interest can be written as
@@ 3.1 Standard g-computation
One approach for estimation of ${\color[rgb]{0,0,0}\mu_{\tau}(\bar{a}_{\tau-1}^{*})}$ is the parametric generalized computation algorithm formula where time-varying variables are simulated forward in time using a series of parametric models [2], which we refer to as ‘standard’ g-computation. Here, we briefly review how standard g-computation with pooled regression models is implemented in order to describe its deficiencies compared to ICE g-computation. More detailed descriptions of how to implement standard g-computation can be found in the following references [5, 4, 7, 28].
@@
As standard g-computation requires one to specify models for $Y$ and each $L$, it is considered to be highly susceptible to bias resulting from model misspecification [4]. Specifically, the standard g-computation estimator is only consistent for ${\color[rgb]{0,0,0}\mu_{\tau}(\bar{a}_{\tau-1}^{*})}$ when each of the time-varying covariate models is correctly specified. This assumption becomes more questionable as the number of time-varying confounding variables, $p$, increases. This is a prominent concern in applied settings, where the number of time-varying covariates is likely to be large. As sufficient information to inform model selection for all time-varying covariates is unlikely in practice, other algorithms for g-computation that do not require one to specify models for $L_{k}$ have been proposed [8].
@@ 3.2 ICE g-computation
${\color[rgb]{0,0,0}\mu_{\tau}(\bar{a}_{\tau-1}^{*})}=\mathbb{E}\left\{...\mathbb{E}\left[\mathbb{E}(Y_{\tau}|\bar{A}_{\tau-1}=\bar{a}_{\tau-1}^{*},\bar{L}_{\tau-1})|\bar{A}_{\tau-2}=\bar{a}_{\tau-2}^{*},\bar{L}_{\tau-2}\right]...|\bar{A}_{0}=\bar{a}_{0}^{*},\bar{L}_{0}\right\}$ (2)
@@
Step 6 provides the point estimate of ${\color[rgb]{0,0,0}\mu_{\tau}(\bar{a}_{\tau-1}^{*})}$, e.g., ${\color[rgb]{0,0,0}\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})}=n^{-1}\sum_{i=1}^{n}\tilde{Y}_{i,0}^{*}$. Unlike standard g-computation, ICE g-computation only requires the analyst to specify $\tau$ regression models for $Y$. In other words, one does not need to specify models for other covariates included in $L_{k}$. For time-to-event data, the described ICE g-computation algorithm is slightly modified (see Appendix 1).
@@
For consistent variance estimation of ${\color[rgb]{0,0,0}\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})}$, the nonparametric bootstrap has been suggested [7, 8]. While the nonparametric bootstrap provides a consistent variance estimator, it can be computationally demanding, as it requires re-estimating the series of regression models using resampled data. Further, the ICE procedure does not provide estimates of ${\color[rgb]{0,0,0}\mu_{\tau-1}(\bar{a}_{\tau-2}^{*})}$ or the mean at other time-points as by-products of estimating ${\color[rgb]{0,0,0}\mu_{\tau}(\bar{a}_{\tau-1}^{*})}$, unlike standard g-computation. Instead, the whole process, and thus also the bootstrap, must be repeated for each time point of interest. Therefore, we consider an alternative to the bootstrap for estimating the variance of ${\color[rgb]{0,0,0}\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})}$.
3.3 ICE g-computation as an M-estimator
To avoid using the bootstrap to estimate the variance, we express the ICE g-computation estimator as an M-estimator. Let $\boldsymbol{\theta}$ be a $v$-dimensional vector, where $v$ is finite. An M-estimator, $\hat{\boldsymbol{\theta}}$, is the solution to the estimating equation
$\sum_{i=1}^{n}\psi(O_{i};\hat{\boldsymbol{\theta}})=\boldsymbol{0}$
@@ 3.3 ICE g-computation as an M-estimator
$\mathbb{F}_{n}(\hat{\boldsymbol{\theta}})=\frac{1}{n}\sum_{i=1}^{n}\left\{\psi(O_{i};\hat{\boldsymbol{\theta}})\psi(O_{i};\hat{\boldsymbol{\theta}})^{T}\right\}.$
@@
For unstratified ICE g-computation with a repeatedly measured binary outcome modeled using logistic regression, the stacked estimating equations are
$\sum_{i=1}^{n}\psi_{U}(O_{i};\hat{\boldsymbol{\theta}})=\sum_{i=1}^{n}\begin{bmatrix}I(C_{i,\tau}=0)\left\{Y_{i,\tau}-\text{expit}(X_{i,\tau-1}^{T}\hat{\beta}_{\tau-1})\right\}X_{i,\tau-1}\\ I(C_{i,\tau-1}=0)\left\{\tilde{Y}_{i,\tau-1}^{*}-\text{expit}(X_{i,\tau-2}^{T}\hat{\beta}_{\tau-2})\right\}X_{i,\tau-2}\\ \vdots\\ I(C_{i,1}=0)\left\{\tilde{Y}_{i,1}^{*}-\text{expit}(X_{i,0}^{T}\hat{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*}-{\color[rgb]{0,0,0}\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})}\end{bmatrix}=\boldsymbol{0}$ (3)
where $\hat{\boldsymbol{\theta}}=(\hat{\beta}_{\tau-1},\hat{\beta}_{\tau-2},...\hat{\beta}_{0},{\color[rgb]{0,0,0}\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})})$, $\text{expit}(\cdot)$ is the inverse logit function, $X_{i,k}$ is the $i$th row of a design matrix composed of user-specified functions of $\bar{A}_{i,k}$ and $\bar{L}_{i,k}$, $\tilde{Y}_{i,k}=\text{expit}(X_{i,k-1}^{*^{T}}\hat{\beta}_{k-1})$, $X_{i,k}^{*}$ is the $i$th row of the design matrix with $\bar{a}_{i,k}^{*}$ replacing $\bar{A}_{i,k}$, and ${\color[rgb]{0,0,0}\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})}$ is the estimator under the specified plan. The first estimating function is the score function of a logistic regression model for the observed outcome at time $\tau$. The second estimating function is the score of a fractional logistic regression model where the dependent variable is the predicted outcome at $\tau$ under the plan [32]. The subsequent estimating functions are recursive fractional logistic regression models backwards through time until baseline. The final estimating function is for the mean of the outcome at $\tau$ under the plan.
As stated, one can show that $\hat{\boldsymbol{\theta}}$ is a consistent and asymptotically normal estimator of $\boldsymbol{\theta}$ by showing that $\hat{\boldsymbol{\theta}}$ is the solution to a vector of unbiased estimating equations. For the estimating equations corresponding to regression models, a generic element $x_{i,\tau-1}$ in the design matrix $X_{i,\tau-1}$ is considered without a loss of generality. For the first estimating equation in 3, note that it can be rewritten as an expectation conditional on $C_{\tau}=0$. Then it follows that
@@
The primary advantage of stacking the estimating equations is that all successive regression models for the ICE g-computation procedure are jointly solved in 3. When unbiased estimating equations are stacked together, standard M-estimation theory is applicable [33]. In other words, the estimated variance of $\hat{\mu}_{\tau}(\bar{a}^{*}_{\tau-1})$ from the empirical sandwich variance estimator correctly incorporates the uncertainty in estimation of nuisance model parameters. This feature is unlike ICE g-computation implementations that solve for the nuisance model parameters in successive steps. For intuition, notice that the bread matrix is defined as the partial derivatives of $\hat{\boldsymbol{\theta}}$ for the stacked estimating equations. Similarly, the meat matrix is defined as the outer product of the stacked estimating equations. Therefore, the empirical sandwich variance estimator incorporates the uncertainty of parameters that are shared between estimating equations in the corresponding stack.
For stratified ICE g-computation with a repeatedly measured binary outcome modeled using logistic regression, the estimating equations are modified to
$\sum_{i=1}^{n}\psi_{S}(O_{i};\check{\boldsymbol{\theta}})=\sum_{i=1}^{n}\begin{bmatrix}I(C_{i,\tau}=0,\bar{A}_{i,\tau-1}=\bar{a}_{i,\tau-1}^{*})\left\{Y_{i,\tau}-\text{expit}(X_{i,\tau-1}^{T}\check{\beta}_{\tau-1})\right\}X_{i,\tau-1}\\ I(C_{i,\tau-1}=0,\bar{A}_{i,\tau-2}=\bar{a}_{i,\tau-2}^{*})\left\{\tilde{Y}_{i,\tau-1}^{*}-\text{expit}(X_{i,\tau-2}^{T}\check{\beta}_{\tau-2})\right\}X_{i,\tau-2}\\ \vdots\\ I(C_{i,1}=0,\bar{A}_{i,0}=\bar{a}_{i,0}^{*})\left\{\tilde{Y}_{i,1}^{*}-\text{expit}(X_{i,0}^{T}\check{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*}-{\color[rgb]{0,0,0}\check{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})}\end{bmatrix}=\boldsymbol{0}$ (4)
where $\check{\boldsymbol{\theta}}=(\check{\beta}_{\tau-1},\check{\beta}_{\tau-2},...\check{\beta}_{0},{\color[rgb]{0,0,0}\check{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})})$. These estimating equations are nearly identical to Equation 3, except for the addition of being restricted to those who followed the treatment plan up to the corresponding time. Notice that a similar proof of unbiasedness for the stacked estimating equations follows here, with stratified ICE g-computation only relaxing the parametric modeling assumptions being made.
The score functions of the logistic models in Equations 3 and 4 can also be replaced with the score functions of other generalized linear models. As indicated above, inference for ${\color[rgb]{0,0,0}\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})}$ and ${\color[rgb]{0,0,0}\check{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})}$ can then be made using the empirical sandwich variance estimator. Importantly, software is available which automates computation of the point and variance estimates for a given set of stacked estimating equations [19, 20]. Estimating equations for ICE g-computation with time-to-event data, as opposed to repeated measures, are provided in Appendix 1 and the following reference [9].

### v3

@@ Abstract
Iterated conditional expectation (ICE) g-computation is an estimation approach for addressing time-varying confounding for both longitudinal and time-to-event data. Unlike other g-computation implementations, ICE avoids the need to specify models for each time-varying covariate. For variance estimation, previous work has suggested the bootstrap. However, bootstrapping can be computationally intense. Here, we present ICE g-computation as a set of stacked estimating equations. Therefore, the variance for the ICE g-computation estimator can be consistently estimated using the empirical sandwich variance estimator. Performance of the variance estimator was evaluated empirically with a simulation study. The proposed approach is also demonstrated with an illustrative example on the effect of cigarette smoking on the prevalence of hypertension. In the simulation study, the empirical sandwich variance estimator appropriately estimated the variance. When comparing runtimes between the sandwich variance estimator and the bootstrap for the applied example, the sandwich estimator was substantially faster, even when bootstraps were run in parallel. The empirical sandwich variance estimator is a viable option for variance estimation with ICE g-computation.
@@ 1 Introduction
To consistently estimate the variance for ICE g-computation, previous work has suggested the nonparametric bootstrap [7, 8, 10]. However, the bootstrap is computationally demanding, as it requires repeating the analysis using resamples of the data [11, 12]. This computational complexity can limit the scenarios considered by researchers in practical applications (e.g., exploring alternative treatment plans, varying functional form specifications, sensitivity analyses). The computational complexity also makes simulation experiments difficult. For example, some simulation studies forgo estimation of the variance by bootstrap in each iteration [8, 9]. While other simulation studies have used the bootstrap [10], the computational burden may limit the sample sizes, number of iterations, or scenarios considered. An alternative variance estimator is based on the influence curve, but this estimator is not consistent for ICE g-computation [7].
Here, the ICE g-computation estimator is expressed as a set of estimating equations [13], which allows the asymptotic variance of the ICE estimator to be consistently estimated using the empirical sandwich variance estimator [8]. The empirical sandwich variance estimator has also been used for variance estimation with g-computation in settings without time-varying confounding [14, 15, 16, 17]. Here, we illustrate the use of estimating equations for ICE g-computation in settings with time-varying confounding. The proposed approach has the usual benefits of ICE g-computation, in addition to providing a statistically consistent variance estimator that is more computationally efficient than the nonparametric bootstrap. This approach also allows one to easily estimate the variance for transformations of parameters and incorporate additional nuisance models by stacking estimating equations together. Finally, the ICE g-computation estimator can be readily implemented using existing software for generic estimating equations [18, 19].
The structure of the paper is as follows. In section 2, the data and a sufficient set of identification assumptions for longitudinal data structures are reviewed. Section 3 reviews g-computation estimators in the setting of repeated measures and presents ICE g-computation as stacked estimating equations. The proposed implementation is examined with a simulation study in section 4. In section 5, the proposed ICE g-computation procedure is demonstrated in an illustrative example of estimating the effect of cigarette smoking on prevalent hypertension. Finally, section 6 summarizes the key results and notes how an estimating equation approach for causal effect estimation with longitudinal data can be expanded upon in future work.
@@ 2 Observed data and identification
Let $k\in\{1,...,\tau\}$ index discrete follow-up times. The potential outcome at $\tau$ for unit $i$ under the treatment plan (i.e., a defined sequence of treatments) $\bar{a}_{\tau-1}^{*}=(a_{0}^{*},a_{1}^{*},...,a_{\tau-1}^{*})$ is denoted by $Y_{i,\tau}(\bar{a}_{\tau-1}^{*})$. Only deterministic plans of binary treatments (i.e., units are assigned to specific treatments, as opposed to probabilities of treatments) are considered hereafter. A simple example is always treat, where $\bar{a}_{\tau-1}^{*}=(1,1,...,1)$. The parameter of interest is the mean potential outcome at $\tau$ under a given plan (i.e., the joint effect of treatment), $\mu_{\tau}(\bar{a}_{\tau-1}^{*})=\mathbb{E}[Y_{i,\tau}(\bar{a}_{\tau-1}^{*})]$, where $\mathbb{E}[\cdot]$ is the expected value function.
@@
To identify $\mu_{\tau}(\bar{a}_{\tau-1}^{*})$, we proceed following the assumptions provided in Table 1. Comprehensive discussion of these assumptions can be found elsewhere [4, 8, 21, 22, 23, 24, 25]. Briefly, causal consistency provides a connection between the observed outcomes and time-varying covariates with the potential outcomes and covariates under the treatment plans. Causal consistency, as expressed here, implies both no interference between units and variations of treatment are irrelevant [26]. Treatment exchangeability stipulates that the treatment at time $k$ is independent of the potential outcomes conditional on the history of previous treatment and covariates. In other words, $L$ includes all the confounding variables. In the context of deterministic plans, treatment positivity states that there is a non-zero probability of following the plan up to $k$ conditional on having followed the plan up to $k-1$ and time-varying confounding history that occur. Censoring exchangeability specifies that loss to follow-up at time $k$ is non-informative conditional on $\bar{L}_{k},\bar{A}_{k},\bar{Y}_{k}$. Again, censoring positivity ensures that the expression for censoring exchangeability is well-defined. In addition to these assumptions, we further assume no measurement error. Following these additional assumptions, the parameter of interest can be written as
@@ 3.1 Standard g-computation
One approach for estimation of $\mu_{\tau}(\bar{a}_{\tau-1}^{*})$ is the parametric generalized computation algorithm formula where time-varying variables are simulated forward in time using a series of parametric models [2], which we refer to as ‘standard’ g-computation. Here, we briefly review how standard g-computation with pooled regression models is implemented in order to describe its deficiencies compared to ICE g-computation. More detailed descriptions of how to implement standard g-computation can be found in the following references [5, 4, 7, 27].
@@
As standard g-computation requires one to specify models for $Y$ and each $L$, it is considered to be highly susceptible to bias resulting from model misspecification [4]. Specifically, the standard g-computation estimator is only consistent for $\mu_{\tau}(\bar{a}_{\tau-1}^{*})$ when each of the time-varying covariate models is correctly specified. This assumption becomes more questionable as the number of time-varying confounding variables, $p$, increases. This is a prominent concern in applied settings, where the number of time-varying covariates is likely to be large. As sufficient information to inform model selection for all time-varying covariates is unlikely in practice, other algorithms for g-computation that do not require one to specify models for $L_{k}$ have been proposed [8].
@@ 3.2 ICE g-computation
$\mu_{\tau}(\bar{a}_{\tau-1}^{*})=\mathbb{E}\left\{...\mathbb{E}\left[\mathbb{E}(Y_{\tau}|\bar{A}_{\tau-1}=\bar{a}_{\tau-1}^{*},\bar{L}_{\tau-1})|\bar{A}_{\tau-2}=\bar{a}_{\tau-2}^{*},\bar{L}_{\tau-2}\right]...|\bar{A}_{0}=\bar{a}_{0}^{*},\bar{L}_{0}\right\}$ (2)
@@
Step 6 provides the point estimate of $\mu_{\tau}(\bar{a}_{\tau-1}^{*})$, e.g., $\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})=n^{-1}\sum_{i=1}^{n}\tilde{Y}_{i,0}^{*}$. Unlike standard g-computation, ICE g-computation only requires the analyst to specify $\tau$ regression models for $Y$. In other words, one does not need to specify models for other covariates included in $L_{k}$. For time-to-event data, the described ICE g-computation algorithm is slightly modified (see Appendix 1) [9].
@@
For consistent variance estimation of $\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})$, the nonparametric bootstrap has been suggested [7, 8]. While the nonparametric bootstrap provides a consistent variance estimator, it can be computationally demanding, as it requires re-estimating the series of regression models using resampled data. Further, the ICE procedure does not provide estimates of $\mu_{\tau-1}(\bar{a}_{\tau-2}^{*})$ or the mean at other time-points as by-products of estimating $\mu_{\tau}(\bar{a}_{\tau-1}^{*})$, unlike standard g-computation. Instead, the whole process, and thus also the bootstrap, must be repeated for each time point of interest. Therefore, we consider an alternative to the bootstrap for estimating the variance of $\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})$.
3.3 ICE g-computation as Estimating Equations
To avoid using the bootstrap to estimate the variance, we express the ICE g-computation estimator via estimating equations. Let $\boldsymbol{\theta}$ be a $v$-dimensional vector, where $v$ is finite. The estimator $\hat{\boldsymbol{\theta}}$, often referred to as a M-estimator or Z-estimator [13, 30], is the solution to the estimating equation
${\color[rgb]{0,0,0}\frac{1}{n}}\sum_{i=1}^{n}\psi(O_{i};\hat{\boldsymbol{\theta}})=\boldsymbol{0}$
@@ 3.3 ICE g-computation as Estimating Equations
$\mathbb{M}_{n}(\hat{\boldsymbol{\theta}})=\frac{1}{n}\sum_{i=1}^{n}\left\{\psi(O_{i};\hat{\boldsymbol{\theta}})\psi(O_{i};\hat{\boldsymbol{\theta}})^{T}\right\}.$
@@
For unstratified ICE g-computation with a repeatedly measured binary outcome modeled using logistic regression, the estimating functions are
$\psi_{U}(O_{i};\hat{\boldsymbol{\theta}})=\begin{bmatrix}I(C_{i,\tau}=0)\left\{Y_{i,\tau}-\text{expit}(X_{i,\tau-1}^{T}\hat{\beta}_{\tau-1})\right\}X_{i,\tau-1}\\ I(C_{i,\tau-1}=0)\left\{\tilde{Y}_{i,\tau-1}^{*}-\text{expit}(X_{i,\tau-2}^{T}\hat{\beta}_{\tau-2})\right\}X_{i,\tau-2}\\ \vdots\\ I(C_{i,1}=0)\left\{\tilde{Y}_{i,1}^{*}-\text{expit}(X_{i,0}^{T}\hat{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*}-\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})\end{bmatrix}$ (3)
where $\hat{\boldsymbol{\theta}}=(\hat{\beta}_{\tau-1},\hat{\beta}_{\tau-2},...\hat{\beta}_{0},\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*}))$, $\text{expit}(\cdot)$ is the inverse logit function, $X_{i,k}$ is the $i$th row of a design matrix composed of user-specified functions of $\bar{A}_{i,k}$ and $\bar{L}_{i,k}$, $\tilde{Y}_{i,k}=\text{expit}(X_{i,k-1}^{*^{T}}\hat{\beta}_{k-1})$, $X_{i,k}^{*}$ is the $i$th row of the design matrix with $\bar{a}_{i,k}^{*}$ replacing $\bar{A}_{i,k}$, and $\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})$ is the estimator under the specified plan. The first estimating function is the score function of a logistic regression model for the observed outcome at time $\tau$. The second estimating function is the score of a fractional logistic regression model where the dependent variable is the predicted outcome at $\tau$ under the plan [32]. The subsequent estimating functions are recursive fractional logistic regression models backwards through time until baseline. The final estimating function is for the mean of the outcome at $\tau$ under the plan.
Following suitable regularity conditions [31, p. 327-329], one can show that $\hat{\boldsymbol{\theta}}$ is a consistent and asymptotically normal estimator of $\boldsymbol{\theta}$ by showing that $\hat{\boldsymbol{\theta}}$ is the solution to a vector of unbiased estimating equations. For the estimating equations corresponding to regression models, a generic element $x_{i,\tau-1}$ in the design matrix $X_{i,\tau-1}$ is considered without a loss of generality. For the first estimating equation in 3, note that it can be rewritten as an expectation conditional on $C_{\tau}=0$, since the estimating function restricts contributions to observations with $C_{\tau}=0$ via the leading indicator function. Then it follows that
@@
The advantage of stacking the estimating equations is that all successive regression models for the ICE g-computation procedure are jointly solved in 3. When finite-dimension estimating equations are stacked together, standard M-estimation theory is applicable [33]. So, the variance of $\hat{\mu}_{\tau}(\bar{a}^{*}_{\tau-1})$ can be consistently estimated by the empirical sandwich variance estimator when the estimating equations for all nuisance model parameters are stacked together. For intuition, notice that the bread matrix is defined as the partial derivatives of $\hat{\boldsymbol{\theta}}$. Similarly, the meat matrix is defined as the outer product of the stacked estimating equations. Therefore, the empirical sandwich variance estimator carries the uncertainty across estimating functions included in the stack.
For stratified ICE g-computation with a repeatedly measured binary outcome modeled using logistic regression, the estimating functions are instead
$\psi_{S}(O_{i};\check{\boldsymbol{\theta}})=\begin{bmatrix}I(C_{i,\tau}=0,\bar{A}_{i,\tau-1}=\bar{a}_{i,\tau-1}^{*})\left\{Y_{i,\tau}-\text{expit}(X_{i,\tau-1}^{T}\check{\beta}_{\tau-1})\right\}X_{i,\tau-1}\\ I(C_{i,\tau-1}=0,\bar{A}_{i,\tau-2}=\bar{a}_{i,\tau-2}^{*})\left\{\tilde{Y}_{i,\tau-1}^{*}-\text{expit}(X_{i,\tau-2}^{T}\check{\beta}_{\tau-2})\right\}X_{i,\tau-2}\\ \vdots\\ I(C_{i,1}=0,\bar{A}_{i,0}=\bar{a}_{i,0}^{*})\left\{\tilde{Y}_{i,1}^{*}-\text{expit}(X_{i,0}^{T}\check{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*}-\check{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})\end{bmatrix}$ (4)
where $\check{\boldsymbol{\theta}}=(\check{\beta}_{\tau-1},\check{\beta}_{\tau-2},...\check{\beta}_{0},\check{\mu}_{\tau}(\bar{a}_{\tau-1}^{*}))$. These estimating equations are nearly identical to Equation 3, except for the addition of being restricted to those who followed the treatment plan up to the corresponding time. Notice that a similar proof of unbiasedness for the stacked estimating equations follows here, with stratified ICE g-computation only relaxing the parametric modeling assumptions being made.
The score functions of the logistic models in Equations 3 and 4 can also be replaced with the score functions of other generalized linear models. As indicated above, inference for $\hat{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})$ and $\check{\mu}_{\tau}(\bar{a}_{\tau-1}^{*})$ can then be made using the empirical sandwich variance estimator. Importantly, software is available which automates computation of the point and variance estimates for a given set of stacked estimating equations [18, 19].